Given that Ae^×+B^-×=4coshx-5sinhx. Determine the values of A and B

1
Expert's answer
2022-06-02T18:23:03-0400

Solution

From definitions cosh(x) = (ex + e-x)/2, sinh(x) = (ex – e-x)/2

So 4cosh(x) – 5sinh(x) = 2(ex + e-x) – 2.5(ex – e-x) = -0.5 ex + 4.5 e-x = -0.5 ex + eln(4.5)-x = -0.5 ex + e-(x- ln(4.5))

If  Ae×+B = 4cosh(x) – 5sinh(x) = -0.5 ex + 4.5 e-x then A = -0.5,  

B = 4.5 e-x  => e-xln(B) = e-x+ln(4.5)  =>  xln(B) = x – ln(4.5)  =>  ln(B) = 1 – ln(4.5)/x  => B = e1-ln(4.5)/x  = e / 4.51/x  = e 4.5-1/x  =>  B = B(x) = e 4.5-1/x    

But it seeems to me that there is mistake in task equality. May be Ae×+Be = 4cosh(x) – 5sinh(x). In this case A = -0.5, B = 4.5.